The weekly revenue of a business selling gummy bear bags is a continuous function of price.

Weekly revenue is $6,285 when the price per is $1.80 and the decay rate is 12%.

Revenue is changing exponentially.

What will revenue be when the price is $4.90  (hint: solver!)
Weekly Revenue = $ ________ (nearest $10)

How fast revenue be changing when the price is $4.90
Revenue Rate = (decrease/increase) ______________ $/$ unitless (nearest 1)
Answer #1

Revenue = Price * Quantity

Weekly revenue = 6285$

Price = 1.8$

Quantity = (6285/1.8) = 3491.667 = 3492

New price = 4.9$

New Weekly revenue = 4.9*3492 = 17110$

Change in revenue rate = ((17110/6285)-1)*100

Change in revenue rate = 172%

Revenue rate increase by 172%